OpenStudy (anonymous):

solve the quation q(q+4)=0

OpenStudy (anonymous):

q=0,4

jimthompson5910 (jim_thompson5910):

q(q+4)=0 q = 0 or q+4 = 0 q = 0 or q = -4

jimthompson5910 (jim_thompson5910):

Use the zero product property, then solve each equation for q

OpenStudy (anonymous):

q( q+4)=0 either q=0 or q+4=0 Solve both the equations \[q=0\] \[q+4=0\] What would you get for q? @KayKoo2013